I am trying to figure out a system of equations problem:
-2x=-10y-2
-x=8y=5
Here's what I did:
I chose to use the top equation:
-2x=-10y-2
+2 +2
-2x+2= -10y
/10 /10
Y= -2/10X =2
Then, I plugged that answer into the equation:

-x+8(-2/10x+2) =5
I had no idea of how to finish, but this is what I did..........

I started with what was in the parenthesis and multiplied 8 times -2 and got -- -16/10x.
Then I multiplied 8 times the other 2 and got 16
-x + -16/10x + 16= 5
Then, I added the -x to the -16x and got -17/10x + 16= 5
I then divided 17/10 from both sides and got a -.0647058.... for x

I stopped here, because the answer is not making any sense



Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

-2x=-10y-2
-x + 8y=5
------
You had 2 = signs in the 2nd equation so I changed it.
-------------
Divide thru the 1st equation by "2" to get:
-x = -5y - 1
-x = -8y + 5
------
Substitute for -x to get:
-5y-1 = -8y+5
---
3y = 6
y = 2
-----
Substitute for "y" and solve for "x"::
-x = -5y-1
-x = -10-1
x = 11
---------
Hope that helps.
=======================
Cheers,
Stan H.
-------------

Answer by Boreal(15235)

I am trying to figure out a system of equations problem:
-2x=-10y-2
-x=8y=5
Here's what I did:
I chose to use the top equation:
-2x=-10y-2
+2 +2
-2x+2= -10y
/10 /10
Y= -2/10X + 2/10 HERE IS WHERE YOU WENT WRONG. YOU HAD TO DIVIDE EVERYTHING BY -10, not 10, and the 2 had to be divided as well..
-2x=-10y-2
-x=8y=5 This needs to be clearer. There shouldn't be two equal signs.
The first becomes -2x+2=-10y; divide by -10 (1/5)x+(-1/5)=y
-x=8y+5 ? -x=(8/5)x-(8/5)+5;;; -x=(8/5)x +(17/5)
(-13/5) x=(17/5)
x=-(17/13)
OR, is it
-x=8y-5? then it will be (33/13)
OR, is it
-x+8y=5? then it is -x + (8/5)x-(8/5)=(25/5)
(3/5)x=(33/5)
x=11
-22=-10y-2
y=2
Hope that helps. If it doesn't, then rewrite the second equation and resubmit.